A straight forward answer please

Part A: For ing enzyme catalytic process Catalytic step E +P Substrate binding The forward rate constants are k1 and k2 and reverse rate constant is k.1 Express rate of complex [Enzyme.Substrate] formation and complex breakdown and determine the KM.
Answer #1

For the above equation, given that k1 and k-1 are the forward and reverse rate constants respectively. Let us take k2 as the rate constant for the formation of the product P.

Now, the rate of the complex [ES] formation is = k1 [E][S] ............. (1)

and rate of the complex [ES] breakdown is = k-1 [ES] + k2 [P][E]

Again, considering that the amount of enzyme is too small as compared to that of the substrate S, it may be assumed that [P][E] \approx   [ES]

Therefore equation (3), i.e., the rate of complex breakdown is = (k-1 + k2) [ES]...... (2)

From steady state approximation, d[ES]/dt = 0

Therefore, rate of complex [ES] formation is equal to the rate of complex breakdown

i.e., Equating equation (1) and (2) we have

k1 [E][S] = (k-1 + k2) [ES]

or, [E][S] = (k-1 + k2) [ES] / k1

or, [E][S] = KM [ES]; where KM = (k-1 + k2)/k1

Therefore, KM = [E][S] / [ES]....... (3) ; KM is called Michaelis constant